OpenStudy (anonymous):

Simplify the following expression. x − 4y + 5z − 3x − y + 2z = ? 2x − 5y + 7z −2x − 5y − 7z −2x − 5y + 7z −2x −5y + 3z

OpenStudy (cwrw238):

add like terms eg x - 3x = -2x

OpenStudy (anonymous):

huh?

OpenStudy (cwrw238):

iits a subtract but u can look upon it as add x + (-3x) = -2x

OpenStudy (anonymous):

x − 4y + 5z − 3x − y + 2z Re-arrange them and then solve x-3x-4y-y+5z+2z -2x-5y+7z :)

OpenStudy (anonymous):

It's option number C :)

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

Your welcome :)

OpenStudy (anonymous):

@xoya Evaluate: 16 − 10 ÷ 2 + 3 6 8 14 0

OpenStudy (anonymous):

i got 16 is that correct

OpenStudy (cwrw238):

the rule here is to work out the division first then work from left to right

OpenStudy (anonymous):

i mean i got 14 is that correct

OpenStudy (cwrw238):

16 - 10 / 2 + 3 = 16 - 5 - 3 = 11 - 3 = 8

OpenStudy (cwrw238):

following the following rule PEDMAS in order of priority parentheses, exponent, divide multiply add subtract

OpenStudy (anonymous):

isn't it 16 - 5 +3

OpenStudy (cwrw238):

so you see how i did the division first?

OpenStudy (anonymous):

so then 11 + 3 then that will be 14

OpenStudy (cwrw238):

oh sorry u r right yes 14

OpenStudy (anonymous):

Bob's Car Rentals charges a flat rate of $35 per day plus $3.50 per mile. If the number of miles is represented by "m", which algebraic expression represents the overall rate for the day using "m"? 38.50m 35m + 3.50 35 + 3.50m 31.50m

OpenStudy (anonymous):

@xoya

OpenStudy (anonymous):

@cwrw238